# Bridging Classical and Quantum Machine Learning: Knowledge Transfer from Classical to Quantum Neural Networks Using Knowledge Distillation

### v1

@@ 4.6 Implementation Details
In our experimental setup, we employed PyTorch [63] as the framework for implementing neural network models, with the Qiskit QASM Simulator [64] used for conducting quantum simulations. Hyperparameter optimization was used, enabling the selection of optimal values for the optimizer, learning rate, loss function, temperature for KL-Divergence, and lambda ($\lambda$) ( equation 6 ). The Adam optimizer [56] was chosen for its effectiveness in handling noisy problems and sparse gradients. We set the learning rate to 0.001 and limited the training to 10 epochs to prevent overfitting while allowing sufficient model training. The loss function employed was Cross-Entropy Loss [51] for classification, complemented by KL-Divergence for the knowledge distillation process, with a temperature setting of 3 to soften the output distributions. A $\lambda$ value of 0.9 was utilized to weigh the distillation loss in the overall loss function appropriately. For the quantum experiments, we performed 5000 shots for each measurement to approximate the expectation values with high accuracy and interpret them as probability distributions. These hyperparameters were meticulously chosen based on extensive hyperparameter optimization to ensure the best performing models.
4.7 Results
Table 1: Performance of Quantum and Classical Architectures on MNIST dataset variants. The table is divided into Student, Teacher, and Distillation categories to exhibit the direct training, teacher, and knowledge distillation results respectively.
Category Architecture Dataset Accuracy (%) + $\Delta$ for Student
Student
Quantum, 3 qubit
MNIST 0-5
96.56
Quantum, 3 qubit
MNIST 0-7
93.75
Quantum, 4 qubit
MNIST 0-7
93.83
Teacher
Classical, AlexNet
MNIST 0-5
99.72
Classical, AlexNet
MNIST 0-7
99.51
Classical, LeNet
MNIST 0-5
99.50
Classical, LeNet
MNIST 0-7
99.16
Classical, Ensemble
MNIST 0-5
99.73
Classical, Ensemble
MNIST 0-7
99.58
Distillation
Teacher: AlexNet, Student: Quantum, 3 qubit
MNIST 0-5
97.44
0.88%
Teacher: AlexNet, Student: Quantum, 3 qubit
MNIST 0-7
94.92
1.17%
Teacher: AlexNet, Student: Quantum, 4 qubit
MNIST 0-7
94.17
0.34%
Teacher: LeNet, Student: Quantum, 3 qubit MNIST 0-5 97.89 1.33%
Teacher: LeNet, Student: Quantum, 3 qubit
MNIST 0-7
93.75
0.00%
Teacher: LeNet, Student: Quantum, 4 qubit MNIST 0-7 94.33 0.50%
Teacher: Ensemble, Student: Quantum, 3 qubit
MNIST 0-5
97.33
0.77%
Teacher: Ensemble, Student: Quantum, 3 qubit MNIST 0-7 95.42 1.67%
Teacher: Ensemble, Student: Quantum, 4 qubit
MNIST 0-7
94.00
0.17%
Table 2: Performance of Quantum and Classical Architectures on FashionMNIST dataset variants. Similar to MNIST, the results are divided into the three categories to provide insights into the quantum models’ capabilities with and without the influence of knowledge distillation.
Category Architecture Dataset Accuracy (%) + $\Delta$ for Student
Student
Quantum, 3 qubit
FashionMNIST 0-5
65.00
Quantum, 3 qubit
FashionMNIST 0-7
68.00
Quantum, 4 qubit
FashionMNIST 0-7
74.25
Teacher
Classical, AlexNet
FashionMNIST 0-5
94.90
Classical, AlexNet
FashionMNIST 0-7
91.64
Classical, LeNet
FashionMNIST 0-5
91.57
Classical, LeNet
FashionMNIST 0-7
85.97
Classical, Ensemble
FashionMNIST 0-5
95.67
Classical, Ensemble
FashionMNIST 0-7
91.79
Distillation Teacher: AlexNet, Student: Quantum, 3 qubit FashionMNIST 0-5 83.44 18.44%
Teacher: AlexNet, Student: Quantum, 3 qubit
FashionMNIST 0-7
73.42
5.42%
Teacher: AlexNet, Student: Quantum, 4 qubit
FashionMNIST 0-7
75.42
1.17%
Teacher: LeNet, Student: Quantum, 3 qubit
FashionMNIST 0-5
77.00
12.00%
Teacher: LeNet, Student: Quantum, 3 qubit
FashionMNIST 0-7
68.67
0.67%
Teacher: LeNet, Student: Quantum, 4 qubit FashionMNIST 0-7 75.83 1.58%
Teacher: Ensemble, Student: Quantum, 3 qubit
FashionMNIST 0-5
67.89
2.89%
Teacher: Ensemble, Student: Quantum, 3 qubit FashionMNIST 0-7 74.42 6.42%
Teacher: Ensemble, Student: Quantum, 4 qubit
FashionMNIST 0-7
72.50
0.00%
The results of the conducted experiments are presented in the two comprehensive tables  table 1 and  table 2, which detail the performance of various architectures across the filtered MNIST and FashionMNIST datasets. The results are categorized into three distinct groups: Student, Teacher, and Distillation, to demonstrate the effectiveness of knowledge transfer in quantum neural network training. The ‘Student’ category represents the quantum models directly trained on the datasets. The ‘Teacher’ category includes the performance of classical convolutional neural networks, serving as the teacher models to guide the students during distillation. Lastly, the ‘Distillation’ category shows the outcomes of the student quantum models when knowledge is distilled from the classical teacher models, highlighting the enhancement in performance due to the distillation process. The “+ $\Delta$ for Student" represents the positive change in accuracy of the student quantum models upon training by knowledge distillation. The Architecture column shows the kind of model used, and in case of knowledge distillation, the teacher and student models used in each experiment. In case of the Student and Teacher categories, the positive increase for student is not relevant, hence kept blank using ‘-’. As depicted in  figure 6, the loss curves for the distillation training of the two student models for 3 and 4 qubits for the FashionMNIST subsets are presented. Each subfigure illustrates the decreasing trend of training and validation losses over epochs, indicating effective learning and generalization of the models. The consistent decline in loss values across all models signifies the models’ ability to learn and adapt from the distilled knowledge, underlining the potential of quantum neural networks in assimilating complex patterns from classical pre-trained networks. The convergence of these curves demonstrates the learning of the models, indicative of the successful transfer of knowledge from the teacher models to their quantum counterparts.
Figure 6: Loss curves for distillation training of the two best student models for the FashionMNIST dataset, showcasing the efficiency of knowledge transfer. (a) Represents the learning curve of the best performing 3-qubit model with FashionMNIST 0-5 and AlexNet as the teacher model. (b) Represents the best performing 4-qubit model with FashionMNIST 0-7 and LeNet as the teacher model.
5 Findings and Discussion
Figure 7: Bar plot visually showing the improvement of the student quantum models compared to each distillation experiment performed on the (a) MNIST dataset subsets and (b) FashionMNIST dataset subsets. Each plot is divided into three groups for the three student quantum models. The grey colored bar in each group represents the baseline performance of the student models for the subset mentioned. Subsequent bars represent the improvement for each of the teacher models AlexNet, LeNet and Ensemble model.
Based on the experimental results of the previous section, in this section, we present our findings and discuss upon them. Based on  table 1table 2 and figure 7, we can observe the followings: (1) When used as a teacher model in knowledge distillation, the LeNet and ensemble models exhibit superior performance compared to the simplified AlexNet architecture. This is likely because LeNet was explicitly designed for grayscale images like those in the MNIST and FashionMNIST datasets, allowing it to capture features more effectively than the more complex AlexNet, which required simplification for our experiments. (2) We observe that the average improvement in performance on the FashionMNIST dataset is greater than that on the MNIST dataset. This could be due to FashionMNIST’s inherently lower baseline accuracies, which leaves more room for improvement. This disparity also highlights the third observation: (3) FashionMNIST, with its more varied and complex images, poses a greater classification challenge than MNIST. Yet, our knowledge distillation (KD) approach manages to improve the performance of quantum models on FashionMNIST significantly, indicating the robustness and applicability of KD in enhancing models trained on complex datasets. (4) It can also be observed that the 4-qubit student models do not significantly outperform the 3-qubit models. This suggests that increasing the complexity of the quantum model slightly does not necessarily lead to better learning. This is a testament to our main research objective: reducing the need for quantum resources while improving and maintaining robust performance. (5) The ensemble model performs best for the MNIST 0-7 and FashionMNIST 0-7 datasets, which are the most extensive datasets in our experiments. This model’s superior performance is attributed to its ability to integrate diverse features and learning aspects from multiple classical models, thereby enhancing its predictive accuracy. Interestingly, this advantage is more highlighted in larger datasets. In contrast, the performance gains in smaller datasets are less remarkable, suggesting that ensemble models excel when there is a wealth of variation to learn from, while in more limited datasets, the opportunity to leverage their full potential is not as readily available. (6) Lastly, we note that the performance improvements through distillation are consistent across different dataset filtrations. This consistency underscores the potential for KD to be used in various quantum machine learning applications, especially as quantum hardware continues to evolve and allows for training on larger and more diverse datasets.
The findings signify the potential for classical machine learning architectures to substantially enhance the performance of quantum models through knowledge distillation. The compatibility of the teacher model with the data is crucial, as evidenced by the superior performance of LeNet and ensemble models tailored for grayscale image datasets. The advancements in knowledge distillation demonstrated in this study are particularly promising for complex datasets, suggesting that as quantum computational resources grow, the scope of quantum machine learning applications will broaden significantly. The use of established classical models like ResNet [65], VGG [66], and InceptionNet [67] as teachers in quantum environments might open up new possibilities for cross-domain applications, reducing the reliance on the development of large quantum models and leveraging the extensive resources of classical pre-trained models. This research paves the way for future endeavors where quantum models are integrated into everyday tasks, marking a significant step forward in the evolution of quantum computing applications.
6 Conclusion and Future Work
[…]

### v2

@@ 4.1.5 Implementation Details
In our experimental setup, we employed PyTorch [67] as the framework for implementing neural network models, with TorchQuantum library [68] used for conducting quantum simulations. Hyperparameter optimization was used, enabling the selection of optimal values for the optimizer, learning rate, loss function, temperature for KL-Divergence, and alpha ($\alpha$) (Eq. 23). The Adam optimizer [59] was chosen for its effectiveness in handling noisy problems and sparse gradients. We set the learning rate to 0.001, and run all experiments for 1000 epochs with early stopping regularizer with a patience = 10 epochs to prevent overfitting while allowing sufficient model training. The loss function employed was Cross-Entropy Loss [69] for classification, complemented by KL-Divergence for the knowledge distillation process. The temperature $\tau$ and loss weight $\alpha$ was found to be $\tau$ = 2 and $\alpha$ = 0.4. For the quantum experiments, we perform the default 1024 shots set by TorchQuantum for each measurement to approximate the expectation values with high accuracy and interpret them as probability distributions. We perform each experiment five times and report the mean of the results with the standard deviation to make them more statistically meaningful.
4.2 Proposed Methodology
4.2.1 Problem Formulation
Given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ where $\mathbf{x}_{i}\in\mathcal{X}$ and $y_{i}\in\mathcal{Y}$, the task is to train the student QNN to approximate the mapping of the teacher model. The performance of the QNN can be evaluated by a loss function $\mathcal{L}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, which measures the discrepancy between the predictions of the QNN and the ground truth labels. The objective function for the QNN is given in Eq. 15 as:
The loss of the model outputs, $\mathcal{L}$, is calculated using cross-entropy loss [69] according to Eq. 16:
Figure 9: The architecture diagram of the research methodology followed in this research. Two networks, a classical frozen network with CNN and fully connected layers, and a trainable quantum network with data encoder and QNN is used. The input images are normalized and pre-processed respectively for the classical CNN model and the quantum QNN model. Following that, the data is passed through the respective models. Outputs from the CNN are passed through a fully connected layer to get output logits. Measurements on the qubits of the QNN gives expectation values (or quantum logits). The logits from the classical network and the QNN are used to calculate the KL-divergence, which is added to the loss of the quantum network to aid in backpropagation
The methodology followed in this study is visualized in Fig. 9. The figure shows that after following data pre-processing steps for classical and quantum models respectively, the classical data is passed through a classical network containing teacher CNN model(s), whereas the quantum data is passed through student QNN model(s). After that, using the outputs of the models, the KL-Divergence is calculated and added to the loss of the student model to calculate a total loss which is used to improve performance of the models. Adding the KL-divergence acts as a supervisory signal while updating the parameters of the quantum model. In this teacher-student setup, the student QNN thus tries to mimic the outputs of the teacher CNN and knowledge transfer is accomplished from classical to quantum models solely based on the mimicry. We discuss the used methods in each step of our experiments in the following portion.
4.2.2 Data Normalization
Data normalization is a crucial preprocessing step in many machine learning tasks, including classical and quantum neural networks. Normalization ensures that the input features are on a similar scale, which can improve the convergence rate and generalization performance of the models.
In this study, we employ a widely used normalization technique known as standardization or z-score normalization. This method rescales the features to have a mean of 0 and a standard deviation of 1. For a feature vector $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$, the standardized feature vector $\mathbf{z}=(z_{1},z_{2},\ldots,z_{n})$ is obtained as:
$z_{i}=\frac{x_{i}-\mu_{i}}{\sigma_{i}},$ (17)
where $\mu_{i}$ and $\sigma_{i}$ are the mean and standard deviation of the $i$-th feature, respectively, computed across the entire dataset.
The mean $\mu_{i}$ and standard deviation $\sigma_{i}$ are calculated as:
$\mu_{i}=\frac{1}{N}\sum_{j=1}^{N}x_{ij},$ (18)
$\sigma_{i}=\sqrt{\frac{1}{N}\sum_{j=1}^{N}(x_{ij}-\mu_{i})^{2}},$ (19)
where $N$ is the total number of samples in the dataset, and $x_{ij}$ is the $i$-th feature of the $j$-th sample.
Standardization ensures that the features are centered around 0 and have a uniform scale, preventing any single feature from dominating the others due to its original scale. This normalization technique is particularly beneficial when dealing with heterogeneous features or when the input features have different units or ranges.
For image data, which is the focus of this study, normalization is typically performed by rescaling the pixel values to a specific range, such as [0, 1] or [-1, 1]. This can be achieved by dividing the pixel values by the maximum possible value (e.g., 255 for 8-bit images).
The mean and standard deviation used for normalization are computed on the training set and applied consistently to the validation and test sets during inference. This ensures that the normalization parameters are not influenced by the unseen data, maintaining the integrity of the evaluation process.
4.2.3 Classical Data Pre-processing and Quantum Data Encoding
For neural networks, the data pre-processing steps involve converting the raw data into a suitable format for efficient processing. In this study, we work with image datasets, and the following steps are performed:
  1. 1.
    Conversion to Tensors: The raw image data is converted into PyTorch tensors, which are multi-dimensional arrays that facilitate efficient computations on GPUs or other hardware accelerators.
  2. 2.
    Dataset Creation: The tensor data is organized into a PyTorch Dataset, which provides a convenient way to access and iterate over the data samples during training and evaluation.
  3. 3.
    DataLoader Creation: PyTorch DataLoaders are created from the Datasets, enabling efficient batching, shuffling, and parallel loading of data during the training process.
These pre-processing steps ensure that the data is properly formatted and optimized for efficient processing by the classical neural network models. After the pre-processing steps, for QNNs, we use angle amplitude encoding to encode the classical data into quantum states. Amplitude encoding is a suitable choice for QNNs as it directly maps the classical data into the amplitudes of the quantum states. In amplitude encoding, it is possible to encode $2^{Q}$ features into $Q$ qubits. This is particularly useful in the NISQ era since the number of available qubits is very low in the quantum systems. It is advantageous when dealing with high-dimensional data or complex feature spaces, as the exponential state space of quantum systems can provide a compact and expressive representation. Moreover, amplitude encoding naturally accommodates both continuous and discrete data, eliminating the need for discretization or binarization steps that may lead to information loss. By directly encoding data into the amplitudes of quantum states, QNNs can potentially exploit quantum parallelism and quantum interference effects to perform computations more efficiently than their classical counterparts.
4.2.4 Dimensionality Reduction for Quantum Circuits
Quantum circuits operate on qubits, which have a limited number of dimensions. Consequently, when working with high-dimensional classical data, such as images, it is necessary to reduce the dimensionality of the data before encoding it into the quantum state. In this study, we employ a dimensionality reduction technique that utilizes a fully connected neural network with two layers.
Let $Q$ denote the number of qubits in our quantum circuits, which is 4 or 8 in our experiments. The dimensionality of the classical data needs to be reduced to $2^{Q}$ to be encoded into $Q$ qubits. The dimensionality reduction process is as follows:
1. Flattening: The input image is flattened into a one-dimensional vector, effectively converting the two-dimensional image into a high-dimensional vector representation.
2. Fully Connected Layer 1: The flattened vector is fed into the first fully connected layer, which acts as a non-linear transformation, mapping the high-dimensional input to an intermediate representation.
3. Fully Connected Layer 2: The second fully connected layer takes the output of the first layer and reduces the dimensionality to $2^{Q}$, producing a $2^{Q}$-dimensional vector that can be directly encoded into the quantum state using angle encoding.
The use of a fully connected neural network for dimensionality reduction offers several advantages over other techniques, such as center cropping, Principal Component Analysis (PCA), or pooling operations like maximum pooling or average pooling. The fully connected layers can learn an optimized transformation that captures the most relevant features from the input data, tailored to the specific task and dataset. In addition, the non-linear activation functions in the fully connected layers allow the network to model complex, non-linear relationships between the input and output dimensions, which may be beneficial for representing the intricate patterns present in the data. Moreover, by incorporating the dimensionality reduction layers within the quantum neural network architecture, the parameters of these layers can be optimized jointly with the rest of the quantum circuit during the training process, potentially leading to better overall performance.
4.2.5 Integration of Knowledge Distillation
Knowledge Distillation aims to transfer knowledge from the teacher model to the student model by minimizing a distillation loss $\mathcal{L}_{KD}$. The process involves a soft target $\tau$ which controls the smoothness of the teacher’s output probability distribution. The distillation loss can be defined using the Kullback-Leibler (KL) divergence [70] in Eq. 20 as:
where $\sigma$ denotes the softmax function and KL denotes the Kullback-Leibler divergence. The softmax function is given according to Eq. 21 by:
$\sigma(\mathbf{z})_{i}=\frac{e^{z_{i}}}{\sum_{j=1}^{K}e^{z_{j}}}$ (21)
And, the KL-divergence is calculated as per Eq. 22:
$KL(P||Q)=\sum_{i=1}^{K}P(i)\log\left(\frac{P(i)}{Q(i)}\right)$ (22)
where $P$ and $Q$ are discrete probability distributions. $P(i)$ represents the probability of the $i$-th class according to the true label distribution, typically obtained from the softmax function of the teacher model $\sigma(f_{t}(\mathbf{x}_{i}))$, and $Q(i)$ is the probability of the $i$-th class according to the predicted label distribution from the student model $\sigma(f_{s}(\mathbf{x}_{i}))$. $K$ is the total number of classes. The KL divergence measures the difference between two probability distributions over the same variable. The total loss for the student QNN, incorporating both the original objective and the distillation loss, is given in Eq. 23 by:
$\mathcal{O}_{total}(f_{s})=\alpha\mathcal{L}_{KD}(f_{s},f_{t})+(1-\alpha)\mathcal{O}(f_{s}),$ (23)
where $\alpha$ is a hyperparameter that balances the importance of the original loss and the distillation loss. In order to balance the supervision from the distillation loss, the overall loss is balanced as $0\leq\alpha\leq 1$.
The student QNN is trained by optimizing $\mathcal{O}_{total}$ through a quantum optimization algorithm. The goal of the optimization is to minimize the loss, which can be formulated in Eq. 24 as:
$f_{s}^{*}=\arg\min_{f_{s}}\mathcal{O}_{total}(f_{s}).$ (24)
5 Results and Analysis
5.1 Preliminary Results
Figure 10: Performance of the 4-qubit PQC student model with different values of hyperparameters
Initially, we set up experiments to find the best hyperparameters. In the case of knowledge distillation, hyperparameters $\tau$ and $\alpha$ play the most important role in the model performance. Thus, we create the hyperparameter space $\tau$ = [1, 2, 5] and $\alpha$ = [0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0], and perform experiments by dividing the training sets into training and validation sets. This helped us to select the optimal values of temperature ($\tau$) and alpha ($\alpha$), which are crucial for the knowledge distillation process. The weight of the distillation is represented by $\alpha$, while $1-\alpha$ represents the weight of the main training loss. Fig. 10 shows a heatmap of the performance of the student model (4-qubit PQC) with varying $\alpha$ and $\tau$ values, using AlexNet as the teacher model.
From the heatmap, we observe that the student model benefits from smaller values of $T$ and $\alpha$ in general. Specifically, lower temperatures tend to sharpen the probability distribution, which helps the student model learn more confidently from the teacher’s predictions. This signifies that the student model benefits from a more focused and less noisy learning signal. Similarly, a smaller $\alpha$ indicates a lower reliance on the distilled knowledge from the teacher, which seems to improve the student’s performance. This implies that the student model can effectively leverage the teacher’s knowledge to achieve higher accuracy with lower weightage on supervisory signals through the distillation process. Concurrently, it implies that the student model possesses a certain degree of inherent capability or intuition that allows it to perform well even without heavily relying on the distilled knowledge from the teacher. This inherent capability might be attributed to the student model’s structure, its initial conditions, or its ability to generalize from limited information. The observed trend suggests that while knowledge distillation helps in improving the student model’s performance, overly depending on the teacher’s output could potentially constrain the student model’s ability to explore and learn from the data itself.
5.2 Main Results
Table 1: Performance of quantum student models and classical teacher models on the MNIST, FashionMNIST, and CIFAR10 datasets, with and without knowledge distillation. T represents teacher and S represents the student model in distillation experiments.
Dataset Category Architecture Accuracy (%)
MNIST Student 4-qubit $77.69\pm 0.25$
8-qubit $90.40\pm 0.15$
Teacher AlexNet $99.44\pm 0.03$
LeNet $98.99\pm 0.05$
Distillation T: AlexNet, S: 4-qubit $85.18\pm 0.20$ (+7.49)
T: LeNet, S: 4-qubit $84.07\pm 0.18$ (+6.38)
T: AlexNet, S: 8-qubit $91.27\pm 0.12$ (+0.87)
T: LeNet, S: 8-qubit $90.63\pm 0.10$ (+0.23)
FashionMNIST Student 4-qubit $77.36\pm 0.30$
8-qubit $83.44\pm 0.20$
Teacher AlexNet $93.68\pm 0.08$
LeNet $90.00\pm 0.12$
Distillation T: AlexNet, S: 4-qubit $80.61\pm 0.25$ (+3.25)
T: LeNet, S: 4-qubit $80.49\pm 0.22$ (+3.13)
T: AlexNet, S: 8-qubit $85.23\pm 0.18$ (+1.79)
T: LeNet, S: 8-qubit $84.71\pm 0.15$ (+1.27)
CIFAR10 Student 4-qubit $25.44\pm 0.40$
8-qubit $31.62\pm 0.35$
Teacher ResNet50 $83.38\pm 0.15$
DenseNet121 $88.24\pm 0.12$
EfficientNetB0 $87.38\pm 0.10$
Distillation T: ResNet50, S: 4-qubit $29.13\pm 0.35$ (+3.69)
T: DenseNet121, S: 4-qubit $29.05\pm 0.32$ (+3.61)
T: EfficientNetB0, S: 4-qubit $27.32\pm 0.28$ (+1.88)
T: ResNet50, S: 8-qubit $34.04\pm 0.30$ (+2.42)
T: DenseNet121, S: 8-qubit $34.22\pm 0.28$ (+2.60)
T: EfficientNetB0, S: 8-qubit $32.87\pm 0.25$ (+1.25)
We present the experimental results on the three datasets in Table 1. From the results, we can observe several patterns and trends. (i) The knowledge distillation process from classical to quantum models demonstrates substantial gains in accuracy across all distillation experiments. We observe an average accuracy improvement of 3.74% on the MNIST, 2.36% on the FashionMNIST, and 2.58% on the CIFAR10 dataset. This shows that knowledge distillation from classical to quantum neural networks is not only possible, but can be applied across different datasets and models. (ii) The performance of QNN models as students shows a marked improvement when the number of qubits is increased from 4 to 8 across all datasets. This suggests that the capacity of QNNs scales significantly with the number of qubits, which is a promising sign for the scalability of quantum models. (iii) The distillation benefits are more prominent for smaller quantum models. The 4-qubit student models gain more in terms of accuracy percentage points compared to the 8-qubit models when distilled from the same teacher architecture. This could indicate that smaller QNNs are more receptive to the distilled knowledge, possibly due to their lower initial performance providing a greater scope for improvement. (iv) The ResNet model, despite having lower baseline accuracy than DenseNet and EfficientNet, improves the accuracy of the student models consistently more than EfficientNet and also more than DenseNet for the 4-qubit model. This can be attributed to the architectural characteristics of ResNet. ResNet’s design incorporates residual blocks with skip connections that facilitate the flow of gradients during training. This allows the student models to benefit from deeper feature hierarchies without the hindrance of vanishing gradients, which is particularly advantageous for smaller models like the 4-qubit QNNs. In contrast, EfficientNet relies on compound scaling and a balance of network width, depth, and resolution for efficient performance. While this makes EfficientNet highly efficient, it may not provide the same level of rich feature extraction as ResNet, which is crucial for knowledge distillation. DenseNet, with its densely connected architecture, encourages feature reuse and reduces the number of parameters. However, the direct connections between all layers may not be as effective in transferring knowledge to a quantum model due to the potential complexity and noise in the feature maps being transferred. Therefore, the geometric properties of ResNet, which simplify and enhance the representational capacity of the data, make it a more suitable teacher for the student models in the context of knowledge distillation. (v) The teacher models’ performance serves as a high benchmark for the student QNNs. While the student models do not yet match the teachers’ accuracy, the gap narrows significantly post-distillation. This narrowing gap is indicative of the potential for QNNs to approach the performance of classical models, given sufficient advancements in quantum computing and model design in the future.
5.3 Statistical Significance
To validate the improvements in accuracy obtained through knowledge distillation, we conducted paired t-tests to assess the statistical significance of our results. The p-values for the comparisons between student models with and without knowledge distillation are presented in Table 2.
Table 2: P-values for paired t-tests on the performance improvements due to knowledge distillation.
Dataset Model P-value
MNIST 4-qubit $1.11\times 10^{-39}$
8-qubit $3.61\times 10^{-18}$
FashionMNIST 4-qubit $2.89\times 10^{-26}$
8-qubit $2.29\times 10^{-23}$
CIFAR10 4-qubit $6.58\times 10^{-24}$
8-qubit $3.51\times 10^{-20}$
The results demonstrate that the accuracy improvements obtained through knowledge distillation are statistically significant across all datasets and model architectures, with p-values significantly less than 0.05. This confirms the effectiveness of our knowledge distillation approach from classical to quantum neural networks.
5.4 Ablation Studies
In this part, we conduct several ablation studies to understand the contribution and effect of the selected methodologies to the overall performance of the student models. To this goal, we conduct ablation studies on the dimensionality reduction techniques and the data encoding techniques applied to the student model, and show a comparative analysis of the performance of the models, as well as show that the best methods were selected.
5.4.1 Ablation on Dimensionality Reduction Techniques
In Sec. 4.2.4, we mention various dimensionality techniques other than using a fully connected (FC) layer that can be applied to the classical data in order to reduce the dimension of the data and prepare it for quantum processing. These techniques are: (i) Center cropping: which includes cropping the classical images at the center with a dimension of $N\times N$, where $N\times N=2^{Q}$ such that these $2^{Q}$ features can be encoded using amplitude encoding. (ii) Principal Component Analysis (PCA): This is a linear dimensionality reduction technique that aims to find a lower-dimensional subspace that captures the maximum variance in the data. PCA projects the high-dimensional data onto a new set of orthogonal axes (principal components) ordered by their variance. By selecting the top principal components, the dimensionality of the data can be reduced while retaining most of the relevant information. (iii) Maximum Pooling: This is a dimensionality reduction technique commonly used in convolutional neural networks. It partitions the input into non-overlapping regions and selects the maximum value from each region, effectively downsampling the input while retaining the most salient features. (iv) Average Pooling: Similar to maximum pooling, average pooling divides the input into non-overlapping regions but instead of selecting the maximum value, it computes the average value within each region. This technique helps to reduce the spatial dimensionality while preserving essential information. A visual representation of the dimensionality reduction techniques along with the original technique applied is shown in Fig. 11.
Figure 11: The dimensionality reduction techniques used to compress the classical data for quantum processing. (a) shows the approach with fully connected layers. (b) shows the approach with maximum pooling. (c) shows the appraoch with average pooling. (d) shows the approach with center cropping, and (e) shows the approach with PCA
In Table 3, we present the ablation studies of these techniques using the two student models on all three datasets, compared to the original results. Based on the ablation studies, we make the following observations: (i) Applying dimensionality reduction techniques such as center cropping, PCA, maximum pooling, and average pooling leads to a significant drop in the classification accuracy compared to using the original FC layers. This underperformance is observed across all three datasets (MNIST, FashionMNIST, and CIFAR10) for both the 4-qubit and 8-qubit student quantum models. (ii) The original method of using FC layers outperforms all dimensionality reduction techniques by a substantial margin. This is a testament to the number of learnable parameters in the fully-connected layer, which helps in retaining information better than the other methods. For instance, on the MNIST dataset, the 4-qubit model achieves an accuracy of 77.69% with FC layers, while the best dimensionality reduction technique (average pooling) achieves only 45.87%. (iii) Among the dimensionality reduction techniques, average pooling consistently performs better than the other techniques (center cropping, PCA, and maximum pooling) across all datasets and models, suggesting that average pooling is relatively more effective in retaining essential features for quantum processing. (iv) The performance gap between the 4-qubit and 8-qubit models is maintained even when using dimensionality reduction techniques, with the 8-qubit model consistently outperforming the 4-qubit model, indicating that the increased qubit capacity allows for better representation and processing of the input data.
Table 3: Accuracy (%) of the student quantum models with different dimensionality reduction techniques compared to the original method applied on the three datasets.
Model MNIST
Center Crop PCA Max Pool Avg Pool FC layers
4-qubit $31.35\pm 0.80$ $14.98\pm 0.60$ $19.21\pm 0.70$ $45.87\pm 0.90$ 77.69 $\pm$ 0.25
8-qubit $42.72\pm 0.70$ $19.30\pm 0.65$ $25.36\pm 0.75$ $52.50\pm 0.85$ 90.40 $\pm$ 0.15
FashionMNIST
4-qubit $27.91\pm 0.85$ $17.01\pm 0.65$ $21.11\pm 0.75$ $48.29\pm 0.95$ 77.36 $\pm$ 0.30
8-qubit $32.04\pm 0.80$ $18.42\pm 0.70$ $18.57\pm 0.70$ $50.01\pm 0.90$ 83.44 $\pm$ 0.20
CIFAR10
4-qubit $12.14\pm 0.90$ $6.85\pm 0.70$ $8.05\pm 0.80$ $14.32\pm 1.00$ 25.44 $\pm$ 0.40
8-qubit $12.33\pm 0.85$ $7.99\pm 0.75$ $9.83\pm 0.85$ $20.73\pm 0.95$ 31.62 $\pm$ 0.35
5.4.2 Ablation on Data Encoding Techniques
In Sec. 4.2.3, we use amplitude encoding as the encoding method to map the reduced classical data into quantum states, thus enabling us to run the QNN circuits and perform classification. However, in Sec. 3.2, we discuss two other approaches which are widely used in quantum data encoding, especially for QNNs. In Table 4, we present the classification results of the two quantum model on the three datasets using the other two encoding techniques, as compared to amplitude encoding. Based on the ablation studies, we make the following observations: (i) The amplitude encoding method outperforms both angle encoding and qubit encoding techniques for all three datasets (MNIST, FashionMNIST, and CIFAR10) and for both the 4-qubit and 8-qubit student quantum models. (ii) Amplitude and qubit encoding consistently performs better than angle encoding, indicating their effectiveness in preserving relevant information for quantum processing. This can be credited to the fact that both amplitude and qubit encoding has the ability to use $2^{Q}$ features encoded in the quantum states, whereas angle encoding can use only $Q$ features for $Q$ qubits. (iii) The performance gap between the 4-qubit and 8-qubit models is observed across all encoding techniques, with the 8-qubit model achieving higher accuracy than the 4-qubit model, further emphasizing the advantage of increased qubit capacity. (iv) The relative performance differences between the encoding techniques become more pronounced as the dataset complexity increases, with amplitude encoding maintaining a larger advantage over the other techniques on the more challenging CIFAR10 dataset compared to MNIST and FashionMNIST. This indicates that in more complex datasets, there is more room for improvement, and thus, applying better techniques can improve performance significantly.
Table 4: Accuracy (%) of the student quantum models with different data encoding techniques compared to the original amplitude encoding method applied on the three datasets.
Model MNIST
Angle encoding Qubit encoding Amplitude encoding
4-qubit 62.21 $\pm$ 0.17 75.04 $\pm$ 0.10 77.69 $\pm$ 0.25
8-qubit 78.94 $\pm$ 0.30 85.66 $\pm$ 0.08 90.40 $\pm$ 0.15
FashionMNIST
4-qubit 68.42 $\pm$ 0.13 76.22 $\pm$ 0.16 77.36 $\pm$ 0.30
8-qubit 71.60 $\pm$ 0.03 81.19 $\pm$ 0.31 83.44 $\pm$ 0.20
CIFAR10
4-qubit 11.90 $\pm$ 0.45 23.65 $\pm$ 0.02 25.44 $\pm$ 0.40
8-qubit 16.14 $\pm$ 0.10 27.80 $\pm$ 0.34 31.62 $\pm$ 0.35
Fig. 12 shows an stacked bar plot of the errors percentage of all the methods applied in this study along with their ablation studies, in descending order of performance for the quantum models on the three datasets. The plot reveals that the quantum models perform exceptionally well on the MNIST dataset compared to the FashionMNIST dataset in all the experiments except for maximum pooling. Furthermore, it shows how poorly maximum pooling and center cropping performs compared to the other techniques. Average pooling performs much better than maximum pooling on all three datasets consistently, implying that for quantum circuits, the average presence of features in a patch of image data aids in extracting features better than retaining the most prominent features in the patch. All the encoding methods perform well since fully-connected layers were used to reduce the dimensionality of the features in these methods. However, angle encoding performs much better than qubit encoding, although they both encode similar number of input features in quantum states. Hence, it can be concluded that amplitude encoding extracts features better than qubit encoding across datasets and quantum models for the image classification task.
Figure 12: Bar plot visually comparing the performance of the quantum models on all the methods applied along with their ablation studies in ascending order of performance
6 Discussion
In our study, we have introduced a novel approach to enhancing the performance of quantum neural networks (QNNs) through knowledge distillation from classical neural networks. A key aspect of our investigation lies in the stark disparity in the number of parameters between the classical (teacher) and quantum (student) models. This section delves into the implications of this disparity, the efficiency of quantum models, and the role of knowledge distillation in bridging the performance gap.
6.1 Parameter Efficiency and Quantum Advantage
Table 5 presents a comparison of the number of parameters across different models used in our experiments.
Table 5: Comparison of Parameter Counts in Classical and Quantum Models
Model Type Model Name Parameter Count
Quantum (Student) 4-qubit 3,193
Quantum (Student) 8-qubit 6,373
Classical (Teacher) AlexNet 94,672,074
Classical (Teacher) LeNet 44,426
Classical (Teacher) ResNet50 23,528,522
Classical (Teacher) DenseNet121 6,964,106
[…]